OpenStudy (anonymous):

Write u(x) = 1/x^2-4 as a composition of simpler functions. Then determine the domain of u(x)

OpenStudy (mathmale):

Think: f(x) = 1/x and g(x) = x^2-4 Is it possible to form a composite function that uses both of these simpler functions? If so, show how, being careful to use proper labels.

OpenStudy (anonymous):

when we say a composite function, what you mean is f(g(x))?